# Ego-Exo4D: Understanding Skilled Human Activity from First- and Third-Person Perspectives

### v3

@@ Abstract
We present Ego-Exo4D, a diverse, large-scale multimodal multiview video dataset and benchmark challenge. Ego-Exo4D centers around simultaneously-captured egocentric and exocentric video of skilled human activities (e.g., sports, music, dance, bike repair). 740 participants from 13 cities worldwide performed these activities in 123 different natural scene contexts, yielding long-form captures from 1 to 42 minutes each and 1,286 hours of video combined. The multimodal nature of the dataset is unprecedented: the video is accompanied by multichannel audio, eye gaze, 3D point clouds, camera poses, IMU, and multiple paired language descriptions—including a novel “expert commentary" done by coaches and teachers and tailored to the skilled-activity domain. To push the frontier of first-person video understanding of skilled human activity, we also present a suite of benchmark tasks and their annotations, including fine-grained activity understanding, proficiency estimation, cross-view translation, and 3D hand/body pose. All resources are open sourced to fuel new research in the community.
@@ 1 Introduction
We contend that both the egocentric and exocentric viewpoints are critical for capturing human skill. Firstly, the two viewpoints are synergistic. The first-person (ego) perspective captures the details of close-by hand-object interactions and the camera wearer’s attention, whereas the third-person (exo) perspective captures the full body pose and surrounding environment context. See Figure 1. Not coincidentally, instructional or “how-to" videos often alternate between a third-person view of the demonstrator and a close-up view of their near-field demonstration. For example, a chef may describe their approach and the equipment from an exo view, then cut to clips showing their hands manipulating the ingredients and tools from an ego-like view.
@@
Secondly, not only are the ego and exo viewpoints synergistic, but there is a need to translate fluently from one to the other when acquiring skill. For example, imagine watching an expert repair a bike tire, juggle a soccer ball, or fold an origami swan—then mapping their steps to your own body. Cognitive science tells us that even from a very young age we can observe others’ behavior (exo) and map it onto our own (ego) [42, 108], and this actor-observer translation remains the foundation of visual learning.
@@
Realizing this potential, however, is not possible using today’s datasets and learning paradigms. Existing datasets comprised of both ego and exo views (i.e., ego-exo) are few [145, 139, 76, 77, 127], small in scale, lack synchronization across cameras, and/or are too staged or curated to be resilient to the diversity of the real world. Thus the current literature for activity understanding primarily attends to either the ego [28, 47] or exo [67, 48, 105, 149] view, leaving the ability to move fluidly between the first- and third-person perspectives out of reach. Instructional video datasets [103, 159, 207, 204] offer a compelling window into skilled human activity, but (like the above) are limited to single-viewpoint video, whether purely exocentric or mixed with “ego-like" views at certain time points.
@@
We introduce Ego-Exo4D, a foundational dataset to support research on ego-exo video learning and multimodal perception. The result of a two-year effort by a consortium of 15 research institutions, Ego-Exo4D is a first-of-its-kind large-scale multimodal multiview dataset and benchmark suite. It constitutes the largest public dataset of time-synchronized first- and third- person video, captured by 740 diverse camera wearers in 123 distinct scenes and 13 cities worldwide. For every sequence, Ego-Exo4D provides both the camera wearer’s egocentric video, as well as multiple (4-5) exocentric videos from tripods placed around the camera wearer. All views are time-synchronized and precisely localized in a metric, gravity-aligned frame of reference. The total collection has 1,286 hours of video and 5,035 instances, each spanning 1 to 42 min. of continuous capture.
@@
Ego-Exo4D is not only multiview, it is also multimodal. Captured with the unique open-source Aria glasses [38], all ego video is accompanied by 7-channel audio, IMU, eye gaze, both RGB and two grayscale SLAM cameras, and 3D environment point clouds. Additionally, Ego-Exo4D provides multiple new video-language resources, all time indexed: first-person narrations by the camera wearers describing their own actions; third-person play-by-play descriptions of every camera wearer action; and third-person spoken expert commentary critiquing their performance. The latter is particularly novel: performed by domain-specific experienced coaches and teachers, it focuses on how an activity is executed rather than merely what is being done, surfacing subtleties in skilled execution not perceivable by the untrained eye. To our knowledge, there is no prior video resource with such extensive and high quality multimodal data.
Alongside this data, we introduce benchmarks for foundational tasks for ego-exo video. We propose four families of tasks: 1) ego-exo relation, for relating the actions of a teacher (exo) to a learner (ego) by estimating semantic correspondences and translating viewpoints; 2) ego(-exo) recognition, for recognizing fine-grained keysteps and task structure; 3) ego(-exo) proficiency estimation, for inferring how well a person is executing a skill; and 4) ego pose, for recovering skilled 3D body and hand movements from ego-video. We provide annotations for each task—the result of more than 200,000 hours of annotator effort. To kickstart work in these new challenges, we also develop baseline models and report their results (Appendices). We are hosting the first public benchmark challenges in 2024.
@@ 2 Related work
Next we review prior work in datasets, human skill, and cross-view analysis. Section 5 will discuss related work for each benchmark task. Table 2 in Appendix 9 summarizes Ego-Exo4D’s properties vs. existing datasets.
@@
There has been a surge of interest in egocentric video understanding, facilitated by recent ego-video datasets showing unscripted daily-life activity as in Ego4D [47], EPIC-Kitchens [27, 28, 163], UT Ego [78], ADL [119], and KrishnaCam [147], or procedural activities as in EGTea [81], AssistQ [172], Meccano [126], CMU-MMAC [77], and EgoProcel [10]. Unlike any of the above, Ego-Exo4D focuses on multimodal ego and exo capture, and it is focused on the domain of skilled activities.
@@
Most existing multiview datasets focus on static scenes [20, 175, 151, 128, 176] and objects [133, 173], with limited (exo only) multiview human activity [169, 26]. CMU-MMAC [77] and CharadesEgo [145] are early efforts to capture both ego and exo video. CMU-MMAC [77] features 43 participants in mocap suits who cook 5 recipes in a lab kitchen. In CharadesEgo [145], 71 Mechanical Turkers record 34 hours of scripted scenarios (e.g., “type on laptop, then pick up a pillow") from the ego and exo perspectives sequentially, yielding unsynchronized videos with non-exact activity matches. More recent ego-exo efforts focus on specific activities in one or two environments. Assembly101 [139] and H2O [76] provide time-synced ego and exo video at a lab tabletop where people assemble toy cars or manipulate handheld objects, with 53 and 4 participants, and 513 and 5 hours of footage, respectively. Homage [127] provides 30 hours of ego-exo video from 27 participants in 2 homes doing household activities like laundry.
Compared to any of the prior efforts, Ego-Exo4D offers an order of magnitude more participants, diverse locations, and hours of footage (740 participants, 123 unique scenes, 13 cities, 1,286 hours). Importantly, our focus on skilled tasks takes the participants out of the lab or home and into settings like soccer fields, dance studios, rock climbing walls, and bike repair shops. Such activities also yield a wide variety of full body poses and movements within the scene, beyond using objects at a tabletop. This variety means Ego-Exo4D augments existing 3D human body pose datasets [193, 80, 66, 68, 49]. Finally, compared to any prior ego-exo resource, Ego-Exo4D’s suite of modalities and benchmark tasks are novel and will expand the research directions the community can take for egocentric and/or exocentric video understanding.
@@
Analyzing skill and action quality has received limited attention  [120, 12, 113, 34, 35, 194]. Research in instructional or “how-to" videos is facilitated by (largely exo) datasets like HowTo100M [103] and others [159, 207, 204, 11]. Challenges include grounding keysteps [207, 37, 178, 104, 103, 36, 10, 89], procedural planning [22, 15, 196, 167, 201, 143, 71, 17], learning task structure [107, 205, 37, 4, 9, 202], and leveraging noisy narrations  [89, 104, 103]. A portion of Ego-Exo4D is procedural activities, but unlike the above, it offers simultaneous ego-exo capture. The scale and diversity of our data—including its three forms of language descriptions—widen the avenues for skilled activity understanding research.
@@
There is limited prior work on ego-exo cross-view modeling, arguably due to a lack of high-quality synchronized real-world data. Prior work explores matching people between videos [5, 6, 40, 179, 170] and learning view-invariant [7, 144, 141, 184, 185, 182] or ego features [82]. Beyond the specific case of ego-exo, cross-view methods are explored for translation  [130, 131, 157, 134], novel view synthesis [90, 135, 137, 171, 168, 164, 19], and aerial to ground matching [132, 86]. Ego-Exo4D provides a testbed of unprecedented size and variety for cross-view modeling. In addition, our ego-exo relation tasks (cf. Section 5) surface new challenges in novel-view synthesis with widely varying viewpoints.
@@ 3 Ego-Exo4D dataset
Next we introduce the dataset and its scope. Notably, the video capture was a distributed but coordinated effort performed by 12 research labs. We present the common framework, and reserve site-specific details for Appendix 10.

### v4

@@ Abstract
We present Ego-Exo4D, a diverse, large-scale multimodal multiview video dataset and benchmark challenge. Ego-Exo4D centers around simultaneously-captured egocentric and exocentric video of skilled human activities (e.g., sports, music, dance, bike repair). 740 participants from 13 cities worldwide performed these activities in 123 different natural scene contexts, yielding long-form captures from 1 to 42 minutes each and 1,286 hours of video combined. The multimodal nature of the dataset is unprecedented: the video is accompanied by multichannel audio, eye gaze, 3D point clouds, camera poses, IMU, and multiple paired language descriptions—including a novel “expert commentary” done by coaches and teachers and tailored to the skilled-activity domain. To push the frontier of first-person video understanding of skilled human activity, we also present a suite of benchmark tasks and their annotations, including fine-grained activity understanding, proficiency estimation, cross-view translation, and 3D hand/body pose. All resources are open sourced to fuel new research in the community. https://ego-exo4d-data.org/
keywords:
video understanding, first-person video, egocentric, video-language, 3D, body pose
@@ 1 Introduction
We contend that both the egocentric and exocentric viewpoints are critical for capturing human skill. Firstly, the two viewpoints are synergistic. The first-person (ego) perspective captures the details of close-by hand-object interactions and the camera wearer’s attention, whereas the third-person (exo) perspective captures the full body pose and surrounding environment context. See Figure 1. Not coincidentally, instructional or “how-to” videos often alternate between a third-person view of the demonstrator and a close-up view of their near-field demonstration. For example, a chef may describe their approach and the equipment from an exo view, then cut to clips showing their hands manipulating the ingredients and tools from an ego-like view.
@@
Secondly, not only are the ego and exo viewpoints synergistic, but there is a need to translate fluently from one to the other when acquiring skill. For example, imagine watching an expert repair a bike tire, juggle a soccer ball, or fold an origami swan—then mapping their steps to your own body. Cognitive science tells us that even from a very young age we can observe others’ behavior (exo) and map it onto our own (ego) (Flavell et al., 1981, Newcombe, 1989), and this actor-observer translation remains the foundation of visual learning.
@@
Realizing this potential, however, is not possible using today’s datasets and learning paradigms. Existing datasets comprised of both ego and exo views (i.e., ego-exo) are few (Sigurdsson et al., 2018, Sener et al., 2022, Kwon et al., 2021, la Torre et al., 2009, Rai et al., 2021), small in scale, lack synchronization across cameras, and/or are too staged or curated to be resilient to the diversity of the real world. Thus the current literature for activity understanding primarily attends to either the ego (Damen et al., 2021, Grauman et al., 2022) or exo (Kay et al., 2017, Gu et al., 2018, Monfort et al., 2019, Soomro et al., 2012) view, leaving the ability to move fluidly between the first- and third-person perspectives out of reach. Instructional video datasets (Miech et al., 2019, Tang et al., 2020, Zhukov et al., 2019, Zhou et al., 2018) offer a compelling window into skilled human activity, but (like the above) are limited to single-viewpoint video, whether purely exocentric or mixed with “ego-like” views at certain time points.
@@
We introduce Ego-Exo4D, a foundational dataset to support research on ego-exo video learning and multimodal perception. The result of a two-year effort by a consortium of 15 research institutions, Ego-Exo4D is a first-of-its-kind large-scale multimodal multiview dataset and benchmark suite. It constitutes the largest public dataset of time-synchronized first- and third- person video, captured by 740 diverse camera wearers in 123 distinct scenes and 13 cities worldwide. For every sequence, Ego-Exo4D provides both the camera wearer’s egocentric video, as well as multiple (4-5) exocentric videos from tripods placed around the camera wearer. All views are time-synchronized and precisely localized in a metric, gravity-aligned frame of reference. The total collection has 1,286 hours of video and 5,035 instances, each spanning 1 to 42 minutes of continuous capture.
@@
Ego-Exo4D is not only multiview, it is also multimodal. Captured with the unique open-source Aria glasses (Engel et al., 2023), all ego video is accompanied by 7-channel audio, IMU, eye gaze, both RGB and two grayscale SLAM cameras, and 3D environment point clouds. Additionally, Ego-Exo4D provides multiple new video-language resources, all time indexed: first-person narrations by the camera wearers describing their own actions; third-person play-by-play descriptions of every camera wearer action; and third-person spoken expert commentary critiquing their performance. The latter is particularly novel: performed by domain-specific experienced coaches and teachers, it focuses on how an activity is executed rather than merely what is being done, surfacing subtleties in skilled execution not perceivable by the untrained eye. All three language corpora are time-stamped against the video. To our knowledge, there is no prior video resource with such extensive and high quality multimodal data.
Alongside this data, we introduce benchmarks for foundational tasks for ego-exo video and we formalize them with annotations and evaluation protocols to spur the community’s efforts. We propose four families of tasks:
  1. 1.
    ego-exo relation, for relating the actions of a teacher (exo) to a learner (ego) by estimating semantic correspondences and translating viewpoints;
  2. 2.
    ego(-exo) recognition, for recognizing fine-grained keysteps and task structure;
  3. 3.
    ego(-exo) proficiency estimation, for inferring how well a person is executing a skill;
  4. 4.
    ego pose, for recovering skilled 3D body and hand movements from ego-video.
We provide annotations for each task—the result of more than 200,000 hours of annotator effort. To kickstart work in these new challenges, we also develop baseline models and report their results. We are hosting the first public benchmark challenges in 2024.
Figure 2: Overview of the paper and its sections, including the Ego-Exo4D dataset (Sec. 3.13.23.3), the natural language descriptions collected alongside the dataset (Sec. 4), and the benchmark tasks (Sec. 5).
Though we are motivated by skill learning, Ego-Exo4D is poised for even broader influence, beyond the proposed benchmarks. Whereas existing datasets lack activity modeling in real-world 3D contexts (e.g., restricted to mocap suits and/or lab settings) and existing 3D datasets typically focus on static scenes and objects. Ego-Exo4D is a resource for general 3D vision—such as environment reconstruction, camera relocalization, audio-visual mapping, and many others. Similarly, our novel video-language resources will offer many opportunities for grounding of actions and objects, multimodal representation learning, and language generation. Finally, though our tasks prioritize perception from the “ego-only” perspective, the exo component of our data ensures its utility for the more traditional exo viewpoint too, e.g., for activity recognition and body pose estimation.
@@
Figure 2 provides a roadmap for this paper. After reviewing related work (Sec. 2), we describe the dataset—its contents, camera setup, participants, and our approach to collection (Sec. 3)—followed by an overview of its three forms of natural language annotations (Sec. 4). Finally, we introduce the benchmark tasks organized into the four families described above, outlining the motivation, task definitions, metrics, and baseline results for each (Sec. 5).
Multimodal Egocentric Datasets
EGTEA-Gaze (Li et al., 2018) 2018 V,A,G 32 1 7 106 15k - -
MECCANO (Ragusa et al., 2021) 2021 V,D,G 20 2 1 61 - - -
EK100 (Damen et al., 2022) 2022 V,A 37 45 N/A (97;300)* - - -
Ego4D (Grauman et al., 2022) 2022 V,A,3D,S,G,I 931 74 N/A 110$\dagger$ - - -
HoloAssist (Wang et al., 2023) 2024 V,A,D,G,3D,I 222 ? 20 (49;165)* - ? ?
Multiview Datasets
IXMAS (Weinland et al., 2006) 2006 V 10 1 N/A 11 - - -
MEVA (Corona et al., 2021) 2021 V,T,GPS 100 28 N/A 37 - - -
Ego-Exo Datasets
CMU-MMAC (la Torre et al., 2009) 2009 V,A,M,I 43 1 5 - - - -
Charades-Ego (Sigurdsson et al., 2018) 2018 TODO 71 N/A N/A 157 - - -
LEMMA (Jia et al., 2020) 2020 V,D 8 14 15 (24;64)* - - -
HOMAGE (Rai et al., 2021) 2020 V,A,T,B,Ma 27 10 70 453 - - -
H2O (Kwon et al., 2021) 2021 V,D 4 3 N/A 36 - - 0.5M
Assembly101 (Sener et al., 2022, Ohkawa et al., 2023) 2022 TODO 53 1 101 (24;90) - - 0.2M
EgoExoLearn (Huang et al., 2024) 2024 V,A,G,I 136 7 8 (95;254) - - -
EgoExo4D 2024 V,A,I,G,3D,6D,B,Ma 740 123 43 689 2.2M 9.6M 4.4M
[…]
@@ 2 Related work
Next we review prior work in datasets, human skill, and cross-view analysis. Section 5 will discuss additional related work for each benchmark task.
@@
There has been a surge of interest in egocentric video understanding, facilitated by recent ego-video datasets showing unscripted daily-life activity as in Ego4D (Grauman et al., 2022), EPIC-Kitchens (Damen et al., 2018, 2021, Tschernezki et al., 2023), UT Ego (Lee et al., 2012), ADL (Pirsiavash and Ramanan, 2012), and KrishnaCam (Singh et al., 2016), or procedural activities as in EGTea (Li et al., 2018), AssistQ (Wong et al., 2022), Meccano (Ragusa et al., 2021), CMU-MMAC (la Torre et al., 2009), EgoProcel (Bansal et al., 2022), and HoloAssist (Wang et al., 2023). Unlike any of the above, Ego-Exo4D focuses on multimodal ego and exo capture, and it is focused on the domain of skilled activities.
Many members of our Ego-Exo4D team worked together to create Ego4D (Grauman et al., 2022). The two datasets share some properties: both emphasize unscripted data, with long continuous captures in authentic environments with diverse participants, and both offer novel benchmark tasks and video-language annotations. However, whereas Ego4D focuses on daily-life activity from the egocentric view alone, Ego-Exo4D focuses on skilled activity in specific domains, captures both egocentric and exocentric viewpoints, and is significantly more multimodal. Ego-Exo4D’s language annotations are also broader in scope compared to Ego4D, going beyond play-by-play action narrations to also include first-person how-to descriptions and third-person expert commentary about the skilled activities.
@@
Most existing multiview datasets focus on static scenes (Chang et al., 2017, Xia et al., 2018, Straub et al., 2019, Ramakrishnan et al., 2021, Xiao et al., 2013) and objects (Reizenstein et al., 2021, Wu et al., 2015), with limited (exo only) multiview human activity (Weinland et al., 2006, Corona et al., 2021). CMU-MMAC (la Torre et al., 2009) and CharadesEgo (Sigurdsson et al., 2018) are early efforts to capture both ego and exo video. CMU-MMAC (la Torre et al., 2009) features 43 participants in mocap suits who cook 5 recipes in a lab kitchen. In CharadesEgo (Sigurdsson et al., 2018), 71 Mechanical Turkers record 34 hours of scripted scenarios (e.g., “type on laptop, then pick up a pillow”) from the ego and exo perspectives sequentially, yielding unsynchronized videos with non-exact activity matches. More recent ego-exo efforts focus on specific activities in one or two environments. Assembly101 (Sener et al., 2022) and H2O (Kwon et al., 2021) provide time-synced ego and exo video at a lab tabletop where people assemble toy cars or manipulate handheld objects, with 53 and 4 participants, and 513 and 5 hours of footage, respectively. LEMMA (Jia et al., 2020) contains multi-agent, multi-task activities with 15 common daily tasks, performed by 8 individuals in 14 unique kitchens/living rooms. Homage (Rai et al., 2021) provides 30 hours of ego-exo video from 27 participants in 2 homes doing household activities like laundry. EgoExoLearn (Huang et al., 2024) provides 120 hours of egocentric videos emulating the human demonstration following process with exocentric demonstration videos.
Compared to any of the prior efforts, Ego-Exo4D offers an order of magnitude more participants, diverse locations, and hours of footage (740 participants, 123 unique scenes, 13 cities, 1,286 hours). Importantly, our focus on skilled tasks takes the participants out of the lab or home and into settings like soccer fields, dance studios, rock climbing walls, and bike repair shops. Such activities also yield a wide variety of full body poses and movements within the scene, beyond using objects at a tabletop. This variety means Ego-Exo4D augments existing 3D human body pose datasets (Zhang et al., 2022, Li et al., 2023, Joo et al., 2017, Khirodkar et al., 2023, Guzov et al., 2021). Finally, compared to any prior ego-exo resource, Ego-Exo4D’s suite of modalities and benchmark tasks are novel and will expand the research directions the community can take for egocentric and/or exocentric video understanding. Table 1 summarizes Ego-Exo4D’s properties compared to those of existing datasets.
@@
Analyzing skill and action quality has received limited attention  (Pirsiavash et al., 2014, Bertasius et al., 2017, Parmar and Morris, 2019, Doughty et al., 2018, 2019, Zhang et al., 2023). Research in instructional or “how-to” videos is facilitated by (largely exo) datasets like HowTo100M (Miech et al., 2019) and others (Tang et al., 2020, Zhukov et al., 2019, Zhou et al., 2018, Ben-Shabat et al., 2020). Challenges include grounding keysteps (Zhukov et al., 2019, Elhamifar and Huynh, 2020, Xu et al., 2021, Miech et al., 2020, 2019, Dvornik et al., 2022, Bansal et al., 2022, Lin et al., 2022), procedural planning (Chang et al., 2020, Bi et al., 2021, Zhao et al., 2022, Wang et al., 2023, Zhong et al., 2023, Shvetsova et al., 2022, Ko et al., 2022, Cao et al., 2022), learning task structure (Narasimhan et al., 2023, Zhou et al., 2018, Elhamifar and Huynh, 2020, Alayrac et al., 2016, Ashutosh et al., 2023, Zhou et al., 2023), and leveraging noisy narrations  (Lin et al., 2022, Miech et al., 2020, 2019). A portion of Ego-Exo4D is procedural activities, but unlike the above, it offers simultaneous ego-exo capture. The scale and diversity of our data—including its three forms of language descriptions—widen the avenues for skilled activity understanding research.
@@
There is limited prior work on ego-exo cross-view modeling, arguably due to a lack of high-quality synchronized real-world data. Prior work explores matching people between videos (Ardeshir and Borji, 2016, 2018, Fan et al., 2017, Xu et al., 2018, Wen et al., 2021) and learning view-invariant (Ardeshir and Borji, 2018, Sigurdsson et al., 2018, Sermanet et al., 2018, Yu et al., 2019, 2020, Xue and Grauman, 2023) or ego features (Li et al., 2021). Beyond the specific case of ego-exo, cross-view methods are explored for translation  (Regmi and Borji, 2018, 2019, Tang et al., 2019, Ren et al., 2021, Luo et al., 2024, Cheng et al., 2024), novel view synthesis (Liu et al., 2021, Ren and Wang, 2022, Rombach et al., 2021, Wiles et al., 2020, Watson et al., 2022, Tseng et al., 2023, Chan et al., 2023), and aerial to ground matching (Regmi and Shah, 2019, Lin et al., 2015). Ego-Exo4D provides a testbed of unprecedented size and variety for cross-view modeling. In addition, our ego-exo relation tasks (cf. Section 5) surface new challenges in novel-view synthesis with widely varying viewpoints.
@@ 3 Ego-Exo4D dataset
Next we introduce the dataset and its scope. Notably, the video capture was a distributed but coordinated effort performed by 12 research labs who worked together over nearly two years to create Ego-Exo4D. Importantly, our data collection across the sites was a coordinated effort, with common guidelines, scenarios, and camera rigs. In this way, the dataset is cohesive at the same time it is diverse.
In the following, we first introduce the ego-exo camera rig and time synchronization process (Sec. 3.1). Then we overview the domains and activities that compose the dataset (Sec. 3.2), followed by discussion of the participants’ diverse backgrounds and expertise (Sec. 3.3).
Figure 3: The Aria device used for egocentric recordings.